1)solve 9*6t=9/36
2) solve In p=6.0
3) using the formula R-0.17 + 0.44 log N, determine how many choices would result in a reation time of 0.42 seconds. Round your answer to the nearest whole number.

Answer by stanbon(75887)

1)solve 9*6t=9/36
Divide both sides by 9 to get:
6t=1/36
t=216
------------------
2) solve In p=6.0
p=e^6
p=403.429
---------------
3) using the formula R-0.17 + 0.44 log N, determine how many choices would result in a reation time of 0.42 seconds. Round your answer to the nearest whole number.
Where is the equal sign?
Cheers,
Stan H.